OpenStudy (anonymous):

what is the polar form for x^2-2y^2 = 4 ?

OpenStudy (aravindg):

@mary.rojas Can help on this.

OpenStudy (anonymous):

ahh.. okay. :) i'll tag her name. thank you. :)

OpenStudy (anonymous):

@mary.rojas can you help me po?

OpenStudy (mary.rojas):

Really,never even heard of polar form @AravindG

OpenStudy (aravindg):

I guess you need to bring in sines and cosines.

OpenStudy (anonymous):

\[put~x=r~\cos \theta,y=r~\sin \theta \]

OpenStudy (anonymous):

not it is an of hyperbola

OpenStudy (anonymous):

here paramatric cordinate are \[\tan \theta and \]

OpenStudy (anonymous):

\[r^2\cos ^2\theta-2r^2\sin ^2\theta=4,r^2\left( \cos ^2\theta-2\sin ^2\theta \right)=4\] r=?

OpenStudy (anonymous):

not that use sec theta and tan theta